How many moles of water are needed to react with 10.0 mol of o2


1
Expert's answer
2021-09-20T06:01:47-0400

2C12H4Cl6 + 23O2 + 2H2O = 24CO2 + 12HCl

Mole ratio = 23:2

= 2/23 × 10 = 0.87mol